# references.bib

### v1

@@ Abstract
Continuous action spaces in reinforcement learning (RL) are commonly defined as interval sets. While intervals usually reflect the action boundaries for tasks well, they can be challenging for learning because the typically large global action space leads to frequent exploration of irrelevant actions. Yet, little task knowledge can be sufficient to identify significantly smaller state-specific sets of relevant actions. Focusing learning on these relevant actions can significantly improve training efficiency and effectiveness. In this paper, we propose to focus learning on the set of relevant actions and introduce three continuous action masking methods for exactly mapping the action space to the state-dependent set of relevant actions. Thus, our methods ensure that only relevant actions are executed, enhancing the predictability of the RL agent and enabling its use in safety-critical applications. We further derive the implications of the proposed methods on the policy gradient. Using Proximal Policy Optimization (PPO), we evaluate our methods on three control tasks, where the relevant action set is computed based on the system dynamics and a relevant state set. Our experiments show that the three action masking methods achieve higher final rewards and converge faster than the baseline without action masking.
@@ 1 Introduction
RL can solve complex tasks in robotics [Han2023-roboticRLsurvey], games [shao2019survey], and is used for fine-tuning large language models [ouyang2022training]. Yet, training reinforcement learning (RL) agents is often sample-inefficient due to frequent exploration of actions, which are irrelevant to learning a good policy. Examples for irrelevant actions are driving into a wall for autonomous cars, or tipping a drone off-balance. The global action space is typically large in relation to the relevant actions in each state. Therefore, exploring these actions frequently can introduce unnecessary costs, lead to slow convergence, or even prevent the agent from learning a suitable policy.
@@
Action masking mitigates this problem by constraining the exploration to the set of relevant state-specific actions, which can be obtained based on task knowledge. For example, when there is no opponent within reach in video games, attack actions are masked from the action space [Ye2020-MOBAmasking]. Leveraging task knowledge through action masking usually leads to faster convergence and also increases the predictability of the RL agent as the set of relevant actions can provide additional meaning. For instance, if the set of relevant actions is a set of verified safe actions, action masking can be used to provide safety guarantees [Fulton2018, Krasowski2023b.ProvablySafeRLSurvey].
@@
Whenever the relevant action set is easy to compute, action masking can be seamlessly integrated into RL algorithms and is the quasi-standard for discrete action spaces [Zhong_Yang_Zhao_2024, Shakya2023], e.g., in motion planning [Mirchevska2018, Krasowski2020, Xiaofei2022-masking, Rudolf2022, Feng2023], games [Huang2022, Ye2020-MOBAmasking, Hou2023-maskinggames], and power systems [Tabas2022, Lingyu2023]. However, real-world systems operate in continuous space and discretization to a higher-level decision space might prevent learning optimal policies. Furthermore, simulation of real-world systems is compute-intensive, and developing RL agents for them often requires additional real-world training [zhao_sim-to-real_2020]. Thus, sample efficiency is particularly valuable for these applications.
@@
In this work, we propose three action masking methods for continuous action spaces. They employ expressive convex set representations, i.e., polytopes or zonotopes for the relevant action set. Our action masking methods generalize previous work in [Krasowski2023b.ProvablySafeRLSurvey], which is constrained to intervals as relevant action sets. This extends the applicability of continuous action masking to expressive convex relevant action sets, which is especially useful when action dimensions are coupled, e.g., for a thrust-controlled quadrotor. To integrate the relevant action set into RL, we introduce three methods: the generator mask, which exploits the generator representation of a zonotope, the ray mask, which projects an action into the relevant action set based on radial directions, and the distributional mask, which truncates the policy distribution to the relevant action set. In summary our main contributions are:
@@
  • We introduce continuous action masking based on arbitrary convex sets representing the state-dependent relevant action sets;
@@
  • We evaluate our approach on three benchmark environments that demonstrate the application conditions of our continuous action masking approaches.
1.1 Related literature
Action masking is mainly applied for discrete action spaces [Fulton2018, Mirchevska2018, Krasowski2020, Xiaofei2022-masking, Rudolf2022, Feng2023, Huang2022, Ye2020-MOBAmasking, Hou2023-maskinggames, Tabas2022, Lingyu2023, Krasowski2023b.ProvablySafeRLSurvey, Zhong_Yang_Zhao_2024, varricchione2024a-safemaskingLTL]. Huang et al. [Huang2022] derive the implications on policy gradient RL while evaluating the theoretical findings on real-time strategy games. The authors show that masking actions leads to higher training efficiency and scales better with increased action spaces compared to simply penalizing the agent whenever an irrelevant action is selected. Huo et al. [Hou2023-maskinggames] extend [Huang2022] to off-policy RL and observe similar empirical results.
Action masking for discrete action spaces can be categorized by the purpose of the state-dependent relevant action set. Often, the set is obtained by removing irrelevant actions based on task knowledge [Huang2022, Hou2023-maskinggames, Xiaofei2022-masking, Rudolf2022, Feng2023, Lingyu2023]. For example, actions that would exceed the maximum speed of an areal vehicle [Feng2023]. The relevant action sets are then usually manually engineered. In contrast, a recent study [Zhong_Yang_Zhao_2024] identifies redundant actions based on similarity metrics and removes them automatically to obtain the relevant action set. Another common interpretation of the relevant action set is that it only includes safe actions [Mirchevska2018, Krasowski2020, Tabas2022, Fulton2018, varricchione2024a-safemaskingLTL]. Here, the system dynamics are typically used to verify the safety of actions. Safety is either defined as avoiding unsafe areas or complying with logic formulas. In our experiments, we use the former approach to obtain a relevant (safe) action set.
For continuous action spaces, there is only work on utilizing action masking with intervals as relevant action sets [Krasowski2023b.ProvablySafeRLSurvey]. In particular, the proposed continuous action masking uses a safe action interval set as relevant action set and maps the action space to this set with a straightforward re-normalization operation. In this paper, we generalize previous work for more expressive relevant action sets and demonstrate the applicability to three benchmark environments. Note that the ray mask and generator mask are mathematically equivalent to the approach in [Krasowski2023b.ProvablySafeRLSurvey] if the relevant action set is constrained to interval sets.
@@ 2 Preliminaries
As basis for our derivations of the three masking methods, we provide a concise overview of RL with policy gradients. Further, we define the system dynamics of the RL environment as well as set representations, which will be used to compute the relevant action set.
@@ 2.1 Reinforcement learning with policy gradients
Given a Markov-Decision Process $\left(\mathcal{S},\mathcal{A},T,r,\gamma\right)$, the goal of RL is to learn a parameterized policy $\pi_{\theta}(a|s)$ that maximizes the expected reward $\max_{\theta}\mathbb{E}_{\pi_{\theta}}\sum_{t}\gamma^{t}r(s_{t},a_{t})$, where $r\rightarrow:\mathcal{S}\times\mathcal{A}$ is the reward the agent receives and $\gamma$ is the discount factor for future rewards [sutton_reinforcement_2018]. The action set $\mathcal{A}$, and state set $\mathcal{S}$ are observable and continuous, and the state-transition function $T(s,a,s^{\prime})$ is stationary.
For policy gradient algorithms, learning the optimal policy $\pi_{\theta}^{*}(a|s)$ is achieved by updating its parameters $\theta$ with the policy gradient
@@
where $A_{\pi_{\theta}}(a,s)$ is the advantage function, which represents the expected improvement in reward by taking action $a$ in state $s$ compared to the average action taken in that state according to the policy $\pi_{\theta}(a|s)$ [sutton_policy_1999]. An estimation of the advantage $A_{\pi_{\theta}}(a,s)$ is usually approximated by a neural network.
@@ 2.2 System model and set representations
where $s\in\mathcal{S}\subset\mathbb{R}^{n^{\mathcal{S}}}$, $a\in\mathcal{A}\subset\mathbb{R}^{N}$, and $w\in\mathcal{W}\subset\mathbb{R}^{n^{\mathcal{W}}}$ denote the system state, input, and disturbance, respectively. Input and disturbance are piece-wise constant with a sampling interval $\Delta t$. We assume $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{W}$ to be convex.
A convex set representation well-suited for representing the relevant action set are zonotopes due to the efficiency of operations on them. A zonotope $\mathcal{Z}\subset\mathbb{R}^{N}$ with center $c\in\mathbb{R}^{N}$, generator matrix $G\in\mathbb{R}^{N\times P}$, and scaling factors $\beta\in\mathbb{R}^{P}$ is defined as
$\mathcal{Z}=\left\{c+G\beta\,\big{|}\,\|\beta\|_{\infty}\leq 1\right\}=\mathcal{Z}(c,G)=\langle c,G\rangle_{\mathcal{Z}}.$ (3)
The Minkowski addition $\mathcal{Z}_{1}\oplus\mathcal{Z}_{2}$ of two zonotopes $\mathcal{Z}_{1}$, $\mathcal{Z}_{2}$ and the linear map $M\mathcal{Z}_{1}$ of a zonotope $\mathcal{Z}_{1}$ are given by
@@ 3 Continuous action masking
To apply action masking, a relevant action set $\mathcal{A}^{r}(s)\subset\mathbb{R}^{N}$ has to be available. For our continuous action masking methods, we specifically require the following two assumptions:
@@
The relevant action set $\mathcal{A}^{r}(s)$ is convex and the center and boundary points are computable.
@@
The policy $\pi_{\theta}:S\times\mathcal{A}\rightarrow\mathbb{R}_{+}$ of the agent is represented by a parameterized probability distribution $a\sim\pi_{\theta}(a|s)=\mathbb{P}_{\theta}(a;\theta)$.
Common convex set representations that fulfill Assumption 1 are polytopes or zonotopes, which can express a variety of geometric shapes. Our continuous action masking methods transform the policy $\pi_{\theta}(a|s)$, for which the parameters $\theta$ usually specify a neural network, into the relevant policy $\pi_{\theta}^{r}:S\times\mathcal{A}^{r}\rightarrow\mathbb{R}_{+}$ through an abstract mapping function $\mathcal{F}$
$a^{r}\sim\pi_{\theta}^{r}(a^{r}|s)=\mathcal{F}\big{(}\pi_{\theta}(a|s)\big{)},$ (5)
so that $a^{r}\in\mathcal{A}^{r}$ always holds. The policy gradient updates only necessitate the ability to sample from, and compute the gradient of the policy, hence $\pi_{\theta}^{r}(a^{r}|s)$ does not have to be available in closed form.
In the following, we introduce and evaluate three mapping methods: generator mask, ray mask, and distributional mask, as shown in Fig. 1. Note that by ensuring only actions in the relevant action set $\mathcal{A}^{r}(s)$ are selected, all three methods potentially increase the predictability and safety of the RL agent. We derive the effect of each mapping approach on the gradient of the agent’s objective function for the most common stochastic policy RL algorithm; PPO.
@@ 3.1 Ray mask
The ray mask locally compresses the action set $\mathcal{A}$ into $\mathcal{A}^{r}$, which is accomplished by scaling each action $a$ alongside a ray from the center of $\mathcal{A}^{r}$ to the boundary of $\mathcal{A}$ This allows for sets with substantially different shapes, as shown in Figure 1(a). Specifically, the relevant policy distribution is created by transforming the action $a\in\mathcal{A}\subset\mathbb{R}^{N}$, given by $\pi_{\theta}(a|s)$, with a mapping function $g:\mathcal{A}\rightarrow\mathcal{A}^{r}$:
$a^{r}=g(a)=c+\frac{\lambda_{\mathcal{A}}(a)}{\lambda_{\mathcal{A}^{r}}(a)}(a-c).$ (6)
Here, $\lambda_{\mathcal{A}}(a)$, and $\lambda_{\mathcal{A}^{r}}(a)$ denote the distances to the boundary of the relevant action set and action space, respectively, measured from the center of the relevant action set $c$ in the direction of $a$. Note that computing $\lambda_{\mathcal{A}^{r}}(a)$ for a zonotope requires solving a convex optimization problem, as specified in Appendix A.4. Yet, the ray mask is applicable for all convex sets, for which we can compute the center and boundary points. Since $g(a)$ is bijective, we can apply the change of variables formula to compute the relevant policy
$\pi_{\theta}^{r}(a^{r}|s)=\pi_{\theta}\left(g^{-1}(a^{r})|s\right)\left|\text{det}\left(\frac{\mathrm{d}}{\mathrm{d}a^{r}}g^{-1}(a^{r})\right)\right|,$ (7)
where $g^{-1}(a^{r})=a$ is the inverse of $g$ at $a$. In general, there is no closed form of this distribution available. However, for policy gradient-based RL, we only require to sample from the distribution and compute its policy gradient. Samples from $\pi_{\theta}^{r}(a^{r}|s)$ are created by sampling from the original policy $a\sim\pi_{\theta}(a|s)$, which is transformed with $a^{r}=g(a)$. The policy gradient is derived in the following.
@@
The policy gradient of $\pi_{\theta}^{r}(a|s)$ for the ray mask is given by
$\nabla_{\theta}J\left(\pi_{\theta}^{r}(a|s)\right)=\mathbb{E}_{\pi_{\theta}^{r}}\left[\nabla_{\theta}\log\pi_{\theta}(a|s)A_{\pi_{\theta}^{r}}(a^{r},s)\right],$ (8)
where $A_{\pi_{\theta}^{r}}(a^{r},s)$ is the advantage function associated to $\pi_{\theta}^{r}(a|s)$, and evaluated at the action $a^{r}$.
@@
Since the determinant in (7) is independent of $\theta$, the gradient of it w.r.t. $\theta$ is
$\nabla_{\theta}\text{det}\left(\frac{\mathrm{d}}{\mathrm{d}a^{r}}g^{-1}(a^{r})\right)=0,$ (9)
which simplifies the score function of $\pi_{\theta}^{r}(a|s)$ to
$\nabla_{\theta}\log\pi_{\theta}^{r}(a|s)=\nabla_{\theta}\log\pi_{\theta}(a|s).$ (10)
Combining (10) and the general form of the policy gradient for $\pi_{\theta}^{r}(a|s)$, as in (1), results in (8).

### v2

@@ Abstract
Continuous action spaces in reinforcement learning (RL) are commonly defined as multidimensional intervals. While intervals usually reflect the action boundaries for tasks well, they can be challenging for learning because the typically large global action space leads to frequent exploration of irrelevant actions. Yet, little task knowledge can be sufficient to identify significantly smaller state-specific sets of relevant actions. Focusing learning on these relevant actions can significantly improve training efficiency and effectiveness. In this paper, we propose to focus learning on the set of relevant actions and introduce three continuous action masking methods for exactly mapping the action space to the state-dependent set of relevant actions. Thus, our methods ensure that only relevant actions are executed, enhancing the predictability of the RL agent and enabling its use in safety-critical applications. We further derive the implications of the proposed methods on the policy gradient. Using proximal policy optimization (PPO), we evaluate our methods on four control tasks, where the relevant action set is computed based on the system dynamics and a relevant state set. Our experiments show that the three action masking methods achieve higher final rewards and converge faster than the baseline without action masking.
@@ 1 Introduction
RL can solve complex tasks in areas such as robotics [Han2023-roboticRLsurvey], games [shao2019survey], and large language models [ouyang2022training]. Yet, training reinforcement learning (RL) agents is often sample-inefficient due to frequent exploration of actions, which are irrelevant to learning a good policy. Irrelevant actions are actions that are either physically impossible, forbidden due to some formal specification, or evidently counterproductive for solving the task. Since the global action space is typically large in relation to the relevant actions in each state, exploring these actions frequently can introduce unnecessary costs, lead to slow convergence, or even prevent the agent from learning a suitable policy.
@@
Action masking mitigates this problem by constraining the exploration to the set of relevant state-specific actions, which can be obtained based on task knowledge. For example, when there is no opponent within reach in video games, attack actions are masked from the action space [Ye2020-MOBAmasking]. Leveraging task knowledge through action masking usually leads to faster convergence and also improves the predictability of the RL agent, especially when the set of relevant actions has a specific notion, such as being the set of safe actions. For instance, if the set of relevant actions is a set of verified safe actions, action masking can be used to provide safety guarantees [Fulton2018, Krasowski2023b.ProvablySafeRLSurvey].
@@
When the relevant action set is easy to compute, action masking usually benefits RL by improving the sample efficiency and is the quasi-standard for discrete action spaces [Shakya2023, Zhong_Yang_Zhao_2024], e.g., in motion planning [Feng2023, Krasowski2020, Mirchevska2018, Rudolf2022, Xiaofei2022-masking], games [Hou2023-maskinggames, Huang2022, Ye2020-MOBAmasking], and power systems [Lingyu2023, Tabas2022]. However, real-world systems operate in continuous space and discretizing it might prevent learning optimal policies. Furthermore, simulation of real-world systems is computationally expensive, and developing RL agents for them often requires additional real-world training [zhao_sim-to-real_2020]. Thus, sample efficiency is particularly valuable for these applications.
@@
In this work, we propose three action masking methods for continuous action spaces. They can employ convex set representations, e.g., polytopes or zonotopes, for the relevant action set. Our action masking methods generalize previous work in [Krasowski2023b.ProvablySafeRLSurvey], which is constrained to intervals as relevant action sets. This extends the applicability of continuous action masking to expressive convex relevant action sets, which is especially useful when action dimensions are coupled, e.g., for a thrust-controlled quadrotor. To integrate the relevant action set into RL, we introduce three methods: the generator mask, which exploits the generator representation of a zonotope, the ray mask, which projects an action into the relevant action set based on radial directions, and the distributional mask, which truncates the policy distribution to the relevant action set. In summary, our main contributions are:
@@
  • We introduce continuous action masking based on convex sets representing the state-dependent relevant action sets;
@@
  • We evaluate our approach on four benchmark environments that demonstrate the applicability of our continuous action masking approaches.
2 Related literature
Action masking has been mainly applied to discrete action spaces [Feng2023, Fulton2018, Hou2023-maskinggames, Huang2022, Krasowski2023b.ProvablySafeRLSurvey, Krasowski2020, Lingyu2023, Mirchevska2018, Rudolf2022, Tabas2022, varricchione2024a-safemaskingLTL, Xiaofei2022-masking, Ye2020-MOBAmasking, Zhong_Yang_Zhao_2024]. Huang et al. [Huang2022] derive implications on policy gradient RL and evaluate their theoretical findings on real-time strategy games. They show that masking actions leads to higher training efficiency and scales better with an increasing number of actions than penalizing the agent for selecting irrelevant actions. Huo et al. [Hou2023-maskinggames] have extended [Huang2022] to off-policy RL and have observed similar empirical results.
Action masking for discrete action spaces can be categorized by the purpose of the state-dependent relevant action set. Often, the set is obtained by removing irrelevant actions based on task knowledge [Feng2023, Hou2023-maskinggames, Huang2022, Lingyu2023, Rudolf2022, Xiaofei2022-masking], e.g., executing a harvesting action before goods are produced [Huang2022]. While the relevant action sets are usually manually engineered, a recent study [Zhong_Yang_Zhao_2024] takes a data-driven approach and identifies redundant actions based on similarity metrics. Another common interpretation of the relevant action set is that it only includes safe actions [Fulton2018, Krasowski2020, Mirchevska2018, Tabas2022, varricchione2024a-safemaskingLTL]. These works typically use the system dynamics to verify the safety of actions. A safe action avoids defined unsafe areas or complies with logic formulas. In our experiments, the relevant action sets are either state-dependent safe action sets or a global relevant action set modeling power supply constraints.
For continuous action spaces, there is work on utilizing action masking with multidimensional intervals (hereafter only referred to as intervals) as relevant action sets [Krasowski2023b.ProvablySafeRLSurvey]. In particular, the proposed continuous action masking represents relevant action sets by intervals that reflect safety constraints and employs straightforward re-normalization to map the action space to the relevant action set. In this paper, we generalize to more expressive relevant action sets and demonstrate the applicability to four benchmark environments.
@@ 3 Preliminaries
As the basis for our derivations of the three masking methods, we provide a concise overview of RL with policy gradients. Further, we define the considered system dynamics as well as the set representations used in this work.
@@ 3.1 Reinforcement learning with policy gradients
A Markov decision process is a tuple $\left(\mathcal{S},\mathcal{A},T,r,\gamma\right)$ consisting of the following elements: the observable and continuous state set $\mathcal{S}\subset\mathbb{R}^{n^{S}}$, the action set $\mathcal{A}\subset\mathbb{R}^{N}$, the state-transition distribution $T(s^{\prime}|a,s)$, which is stationary and describes the transition probability to the next state $s^{\prime}\in\mathcal{S}\subset\mathbb{R}^{n^{\mathcal{S}}}$ from the current state $s\in\mathcal{S}$ when executing action $a\in\mathcal{A}$, the reward $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and the discount factor $\gamma$ for future rewards [sutton_reinforcement_2018]. The goal of RL is to learn a parameterized policy $\pi_{\theta}(a|s)$ that maximizes the expected reward $\max_{\theta}\mathbb{E}_{\pi_{\theta}}\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})$.
For policy gradient algorithms, learning the optimal policy $\pi_{\theta}^{*}(a|s)$ is achieved by updating its parameters $\theta$ using the policy gradient [sutton_policy_1999, Thm. 2]
@@
where $A_{\pi_{\theta}}(a,s)$ is the advantage function, which represents the expected improvement in reward by taking action $a$ in state $s$ compared to the average action taken in that state according to the policy $\pi_{\theta}(a|s)$. An estimation of the advantage $A_{\pi_{\theta}}(a,s)$ is usually provided by a neural network.
@@ 3.2 System model and set representations
where $w\in\mathcal{W}\subset\mathbb{R}^{n^{\mathcal{W}}}$ denotes a disturbance. The input is piece-wise constant with a sampling interval $\Delta t$. We assume $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{W}$ to be convex.
Zonotopes are a convex set representation well-suited for representing the relevant action set, due to the efficiency of computing their Minkowski sums and linear maps. A zonotope $\mathcal{Z}\subset\mathbb{R}^{N}$ with center $c\in\mathbb{R}^{N}$, generator matrix $G\in\mathbb{R}^{N\times P}$, and scaling factors $\beta\in\mathbb{R}^{P}$ is defined as
$\mathcal{Z}=\left\{c+G\beta\,\big{|}\,\|\beta\|_{\infty}\leq 1\right\}=\langle c,G\rangle_{\mathcal{Z}}.$ (3)
Additionally, let us denote that $G_{(\cdot,i)}$ returns the $i$-th column vector of the generator matrix $G$. The Minkowski addition $\mathcal{Z}_{1}\oplus\mathcal{Z}_{2}$ of two zonotopes $\mathcal{Z}_{1}$, $\mathcal{Z}_{2}$ and the linear map $M\mathcal{Z}_{1}$ of a zonotope $\mathcal{Z}_{1}$ are given by [althoff2010dissertation, Eq. 2.1]
@@ 4 Continuous action masking
To apply action masking, a relevant action set $\mathcal{A}^{r}(s)\subseteq\mathcal{A}$ has to be available, which constrains the action space $\mathcal{A}$ based on task knowledge. Let us denote the state-dependent relevant action set as $\mathcal{A}^{r}(s)\subseteq\mathcal{A}$. From now on, we omit the dependency on the state to simplify notation. For our continuous action masking methods, we specifically require the following two assumptions:
@@
The relevant action set $\mathcal{A}^{r}$ is convex and its center and boundary points are computable.
@@
The policy $\pi_{\theta}:S\times\mathcal{A}\rightarrow\mathbb{R}_{+}$ of the agent is represented by a parameterized probability distribution $a\sim\pi_{\theta}(a|s)$.
Common convex set representations that fulfill Assumption 1 are polytopes or zonotopes. Our continuous action masking methods transform the policy $\pi_{\theta}(a|s)$, for which the parameters $\theta$ usually specify a neural network, into the relevant policy $\pi_{\theta}^{r}:S\times\mathcal{A}^{r}\rightarrow\mathbb{R}_{+}$ through a functional $h:(\Pi\times\mathcal{P}(\mathcal{A}))\rightarrow\Pi^{r}$. Here, $\Pi$ is the space of all policies, $\Pi^{r}$ is the space of all relevant policies, and $\mathcal{P}(\mathcal{A})$ is the power set of all $\mathcal{A}^{r}$. The transformation is defined as
$a^{r}\sim\pi_{\theta}^{r}(a^{r}|s)=h\big{(}\pi_{\theta}(a|s),\mathcal{A}^{r}\big{)},$ (5)
and thereby ensures that $a^{r}\in\mathcal{A}^{r}$ always holds. Please note that policy gradient methods only require the ability to sample from and to compute the gradient of the policy distribution. Therefore, explicit closed-form expressions for $\pi_{\theta}^{r}(a^{r}|s)$ and $h$ are not necessary.
In the following subsections, we introduce and evaluate three masking methods: generator mask, ray mask, and distributional mask, as shown in Fig. 1. We derive the effect of each masking approach on the gradient of the objective function for stochastic policy gradient methods.
@@ 4.1 Ray mask
The ray mask maps the action set $\mathcal{A}$ to $\mathcal{A}^{r}$ by scaling each action $a$ alongside a ray from the center of $\mathcal{A}^{r}$ to the boundary of $\mathcal{A}$, as shown in Figure 1(a). Specifically, the relevant policy $\pi_{\theta}^{r}$ results from mapping the action $a$, sampled from $\pi_{\theta}(a|s)$, using the function $g_{\mathrm{R}}:\mathcal{A}\rightarrow\mathcal{A}^{r}$:
$a^{r}=g_{\mathrm{R}}(a)=c+\frac{\lambda_{\mathcal{A}^{r}}(a)}{\lambda_{\mathcal{A}}(a)}(a-c).$ (6)
Here, $\lambda_{\mathcal{A}}(a)$ and $\lambda_{\mathcal{A}^{r}}(a)$ denote the distances of the action $a$ to the boundaries of the relevant action set and action space, respectively, measured from the center of the relevant action set $c$ in the direction of $a$. Note that computing $\lambda_{\mathcal{A}^{r}}(a)$ for a zonotope requires solving a convex optimization problem, as specified in Appendix A.4. Yet, the ray mask is applicable for all convex sets, for which we can compute the center and boundary points. Since $g_{\mathrm{R}}(a)$ is bijective (see Appendix A.1 for a detailed proof), we can apply a change of variables [bishop_pattern_2006, Eq. 1.27] to compute the relevant policy
$\pi_{\theta}^{r}(a^{r}|s)=\pi_{\theta}\left(g_{\mathrm{R}}^{-1}(a^{r})|s\right)\left|\text{det}\left(\frac{\mathrm{d}}{\mathrm{d}a^{r}}g_{\mathrm{R}}^{-1}(a^{r})\right)\right|,$ (7)
where $g_{\mathrm{R}}^{-1}(a^{r})=a$ is the inverse of $g_{\mathrm{R}}$. In general, there is no closed form of the distribution $\pi_{\theta}^{r}$ available. However, for stochastic policy gradient-based RL, we only require to sample from $\pi_{\theta}^{r}$ and compute its policy gradient. Samples from $\pi_{\theta}^{r}(a^{r}|s)$ are created by sampling from the original policy $a\sim\pi_{\theta}(a|s)$ followed by computing $a^{r}=g_{\mathrm{R}}(a)$. The policy gradient is derived next.
@@
The policy gradient of $\pi_{\theta}^{r}(a^{r}|s)$ for the ray mask is
$\nabla_{\theta}J\left(\pi_{\theta}^{r}(a^{r}|s)\right)=\mathbb{E}_{\pi_{\theta}^{r}}\left[\nabla_{\theta}\log\pi_{\theta}(a|s)A_{\pi_{\theta}^{r}}(a^{r},s)\right],$ (8)
where $A_{\pi_{\theta}^{r}}(a^{r},s)$ is the advantage function associated with $\pi_{\theta}^{r}(a^{r}|s)$.
@@
The determinant in (7) is independent of $\theta$, i.e.,
$\nabla_{\theta}\text{det}\left(\frac{\mathrm{d}}{\mathrm{d}a^{r}}g_{\mathrm{R}}^{-1}(a^{r})\right)=0,$ (9)
which simplifies the score function of $\pi_{\theta}^{r}(a^{r}|s)$ to
$\nabla_{\theta}\log\pi_{\theta}^{r}(a^{r}|s)=\nabla_{\theta}\log\pi_{\theta}(a|s).$ (10)
Combining (10) and the general form of the policy gradient in (1) for $\pi_{\theta}^{r}(a^{r}|s)$ results in (8).